OpenStudy (anonymous):

Please help!! Will fan and medal!!!! (Am typing the question in the comments it may take a second)

OpenStudy (anonymous):

Find the least common denominator for these two rational expressions. \[\frac{ y }{ y^2-9 } \] & \[\frac{ -3y }{ y^2-4y+3 }\]

OpenStudy (anonymous):

y/(y-3)

OpenStudy (sparrow2):

i think it 1/(y-3) if you ask for a least

OpenStudy (sparrow2):

but not totally sure :(

OpenStudy (anonymous):

u have a common factor in numerator too

OpenStudy (anonymous):

so it will be y/(y-3)

OpenStudy (sparrow2):

yeah but they ask for a least one

OpenStudy (anonymous):

of rational expression

OpenStudy (anonymous):

so should I go with y? cause I thought that 1=y ...

OpenStudy (sparrow2):

okay try with y

OpenStudy (sparrow2):

so do you have other question,mayby i will be more helpful in them :)

OpenStudy (sparrow2):

@welshfella take a look. they ask for least common denominator .and i'm stuck wich to choose: y/(y-3) or 1/(y-3)

OpenStudy (sparrow2):

@welshfella

OpenStudy (anonymous):

Sorry my computer keeps freezing. I'm running a test now and the gonna fix it hopefully. Will you still be on around 3:40-4:00 ?

OpenStudy (welshfella):

y^2 - 9 = (y + 3)(y - 3) y^2 - 4y + 3 = (y - 3)(y - 1) so LCM = ?

OpenStudy (welshfella):

gotta go

OpenStudy (dani_rose):

Hopefully that helps you @WhatEven

OpenStudy (dani_rose):

Videos kinda boring :/ sorry in advance

OpenStudy (welshfella):

Hint: (y -3) is common to both factors

OpenStudy (welshfella):

If there are duplicates only 1 is used to get the LCM

OpenStudy (welshfella):

eg find LCM of 4 and 6 4 = 2*2 6 = 2*3 there are duplicate 2's LCM = 2*2*3

OpenStudy (welshfella):

talking to myself again....